two pipes togather can fill a large tank in 10hrs. one of the pipes , used alone , takes 15hrs longer than the other to fill the tank . hown long would each pipe take to fill the tank alone?
Answer by stanbon(75887)

two pipes togeather can fill a large tank in 10hrs
---------.
Together DATA:
Time = 10 hrs/job ; Rate = 1/10 job/hr
--------------------
one of the pipes , used alone , takes 15hrs longer than the other to fill the tank .
1st pipe DATA;
Time = x hrs/job ; Rate = 1/x job/hr
------------
2nd pipe DATA:
Time = x+15 hrs/job ; Rate = 1/(x+15) job/hr
-------------------
how long would each pipe take to fill the tank alone?
---------------------
EQUATION:
rate + rate = together rate
1/x + 1/(x+15) = 1/10
Multiply thru by 10x(x+15) to get:
10(x+15) + 10x = x(x+15)
20x + 150 = x^2+15x
x^2-5x+150=0
(x-15)(x+10) = 0
Positive solution:
x = 15 hrs (time required for 1st pipe to do the job alone)
x+15 = 30 hrs (time required for the 2nd pipe to do the job alone)
============
Cheers,
Stan H.